unicorn123:

What is the best approximation of the area of a circle with a diameter of 17 meters? Use 3.14 to approximate pi. 53.4 m² 106.8 m² 226.9 m² 907.5 m²

unicorn123:

need help please

MAGA:

what do u think it is?

unicorn123:

maybe 53.4

MAGA:

and how did u get that?

MAGA:

wouldn't it be like 3.14x17. Then what would that be?

unicorn123:

i did that got 53.38 and my closes anser choose is 53.4

MAGA:

yes. Exactly good job

unicorn123:

thanks

MAGA:

@unicorn123 wrote:
thanks
np

unicorn123:

need help What is the best approximation for the area of this circle? Use 3.14 to approximate pi. 12.6 m² 25.1 m² 50.2 m² 158.0 m²

MAGA:

@unicorn123 wrote:
need help What is the best approximation for the area of this circle? Use 3.14 to approximate pi. 12.6 m² 25.1 m² 50.2 m² 158.0 m²
need more info

MAGA:

which one do u think it is?

unicorn123:

i did 3.14 times 4 so i think it is 12.56 i got for the times but the closes anser choose is 12.6m2

carmelle:

The formula for the area of a circle is: \[A = \pi r ^{2}\] The diameter is given, 17. The radius is half of the diameter. So you would do: \[\frac{ 17 }{ 2 }\] which is equal to `8.5`. Now, you can multiply and find the area. \[A = (3.14)(8.5^{2})\] \[A = (3.14)(72.25)\] \[A = ~?\]

Nior:

A=\[πr ^{2}\]=π·172≈907.92028